# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ IV A General Construction for Non-Uniform Fractionalisation
IV.1 Ground State Degeneracy
Clearly, our system has some degree of ground state degeneracy. The only requirement to minimise the energy is that every $\tilde{\theta}$ sits at an integer value, and each field has an infinite number of such values that it can take. Thus, one might naively expect that the ground state has infinite degeneracy. To avoid this contradiction, some configurations of $\tilde{\theta}$ must be equivalent. To see how, note that by bosonising we introduced a spurious degree of freedom. A fermionic field is unchanged under a trivial transformation $\psi\rightarrow e^{2i\pi}\psi$, so adding a multiple of $2\pi$ to the bare fermion bosonic fields also cannot correspond to a physical change in the system. Note that a similar calculation of the ground state degeneracy for a uniformly spaced set of wires, can be found in [170].
With this in mind, let us consider a system of $N$ coupled wires with periodic boundaries in both the $x$ and $y$ directions. That is, the $(N/2-1)^{\textup{th}}$ wire is coupled to the $(-N/2)^{\textup{th}}$ wire. We shall assume that the system is in a ground state, each pair of wires has a uniform value of $\braket{\theta_{j+\frac{1}{2}}(x)}$ pinned to a multiple of $\pi$.
Within this system, we focus (without loss of generality) on a set of three wires at positions $j=0,1,2$, each set to fractions $ua_{j}^{2}$, and introduce the following gauge transformation on the $1^{\textup{st}}$ wire,
$\phi_{1}^{R}$ $\rightarrow\phi_{1}^{R}+2\pi\alpha,$ (59)
$\phi_{1}^{L}$ $\rightarrow\phi_{1}^{L}+2\pi\beta,$ (60)
with parameters $\alpha,\beta\in\mathbb{Z}$. Under this transformation, our sum and difference fields Eqs. 18 and 19 change according to
$\varphi_{1}\rightarrow\varphi_{1}+(\alpha+\beta)\pi,$ (61)
$\theta_{1}\rightarrow\theta_{1}+(\alpha-\beta)\pi.$ (62)
Thus, let us define a new pair of variables $\gamma=\alpha+\beta$ and $\delta=\alpha-\beta$. Given that $\alpha$ and $\beta$ are free to take any integer value, we see that $\gamma$ and $\delta$ can also take any value provided they are both even or both odd. Following the transformation through to the couplings, we find that
$2\tilde{\theta}_{1.5}$ $\rightarrow 2\tilde{\theta}_{1.5}+\left(\gamma+ua_{1}^{2}\delta\right)a_{2}\pi,$ (63)
$2\tilde{\theta}_{0.5}$ $\rightarrow 2\tilde{\theta}_{0.5}+\left(-\gamma+ua_{1}^{2}\delta\right)a_{0}\pi.$ (64)
Let us consider two cases. First we shall look at the case of a uniform Laughlin phase ($a_{j}=1,\forall j$) and verify the known result that a phase with $1/u$ fractionalised quasiparticles should have a $u$-fold degenerate ground state. Then we shall extend the analysis to the general case ($a_{j}\neq 1$), finding the degeneracy of a general composite system.
The Uniform Case: Setting all $a_{j}=1$, we ensure all wires have the same fractionalisation. Let us consider the effect of two gauge transformations, where any other gauge transformation can be produced by successively applying these two. Thus, a quantity that is invariant under these two is gauge invariant in general.
First we look at the transformation parametrised by $\gamma=2,\delta=0$. The effect on our fields is
$\tilde{\theta}_{1.5}$ $\rightarrow\tilde{\theta}_{1.5}+\pi,$ (65)
$\tilde{\theta}_{0.5}$ $\rightarrow\tilde{\theta}_{0.5}-\pi.$ (66)
Successively applying such a transformation on different wires allows one to make any change to the $\tilde{\theta}$ values that preserves $\sum_{j}\tilde{\theta}_{j+\frac{1}{2}}$.
Next, we consider a transformation parametrised by $\gamma=1,\delta=1$. The effect on our fields is
$\tilde{\theta}_{1.5}$ $\rightarrow\tilde{\theta}_{1.5}+\frac{u+1}{2}\pi,$ (67)
$\tilde{\theta}_{0.5}$ $\rightarrow\tilde{\theta}_{0.5}+\frac{u-1}{2}\pi.$ (68)
Thus, we see that under such a transformation, the value of $\sum_{n}\tilde{\theta}_{j+\frac{1}{2}}$ changes by
$\sum_{n}\tilde{\theta}_{j+\frac{1}{2}}\rightarrow\sum_{n}\tilde{\theta}_{j+\frac{1}{2}}+u\pi.$ (69)
Thus, we propose that the only gauge invariant quantity one can derive from the values of $\tilde{\theta}$ is
$A_{GS}=\frac{1}{\pi}\sum_{n}\tilde{\theta}_{j+\frac{1}{2}}\mod u.$ (70)
Since $A_{GS}$ can take $u$ possible integer values, we find that the number of degenerate ground states is
$N_{GS}=u,$ (71)
which is the expected ground state degeneracy for a $1/u$ Laughlin phase.
The Non-uniform Case— We now revisit the same argument, while allowing for $a_{j}\neq 1$. Additionally, let us also require (without loss of generality) that all adjacent $a_{j}$ values are coprime, since any common factors may be absorbed into $u$. First, considering the transformation $\gamma=2,\delta=0$, we find an overall change to our fields of
$\frac{\tilde{\theta}_{1.5}}{a_{1}a_{2}}$ $\rightarrow\frac{\tilde{\theta}_{1.5}}{a_{1}a_{2}}+\frac{\pi}{a_{1}},$ (72)
$\frac{\tilde{\theta}_{0.5}}{a_{0}a_{1}}$ $\rightarrow\frac{\tilde{\theta}_{0.5}}{a_{0}a_{1}}-\frac{\pi}{a_{1}}.$ (73)
Thus, in this case, let us propose the following candidate for our overall conserved quantity:
$A_{GS}=\frac{1}{\pi}\sum_{j}\frac{\tilde{\theta}_{j+\frac{1}{2}}}{a_{j}a_{j+1}}.$ (74)
Next, we once again consider the $\gamma=1,\delta=1$ gauge transformation, which produces a change in our fields of
$\frac{\tilde{\theta}_{1.5}}{a_{1}a_{2}}$ $\rightarrow\frac{\tilde{\theta}_{1.5}}{a_{1}a_{2}}+\left(ua_{1}+\frac{1}{a_{1}}\right)\frac{\pi}{2},$ (75)
$\frac{\tilde{\theta}_{0.5}}{a_{0}a_{1}}$ $\rightarrow\frac{\tilde{\theta}_{0.5}}{a_{0}a_{1}}+\left(ua_{1}-\frac{1}{a_{1}}\right)\frac{\pi}{2},$ (76)
under which the total value of $A_{GS}$ changes by
$A_{GS}\rightarrow A_{GS}+ua_{1}.$ (77)
Similarly, choosing any other wire, we are free to perform such a gauge transformation, adding ±$\pm ua_{j}$ to the value of $A_{GS}$, for any $a_{j}$ present in the system. This means that, starting at an arbitrary value of $A_{GS}$, we may freely gauge transform to any other value that can be written in the form
$A_{GS}\rightarrow A_{GS}+u\sum_{j}w_{j}a_{j},$ (78)
where the sum is over all wires, with arbitrary $w_{j}\in\mathbb{Z}$. Since all $a_{j}$ are coprime, we invoke Bezout’s identity [171], which states that the sum on the right can be made to equal to any multiple of $u$ by choosing appropriate values of $\{a_{j}\}$. Thus, any pair of configurations for which $A_{GS}$ differs by a multiple of $u$ must be gauge equivalent.
All that remains is to determine how many gauge-inequivalent values of $A_{GS}$ exist. Since each $\tilde{\theta}$ field can take any value that is a multiple of $\pi$, the overall value of $A_{GS}$ can always be expressed in the form
$A_{GS}$ $=\frac{1}{\pi}\sum_{j}\frac{\tilde{\theta}_{j+\frac{1}{2}}}{a_{j}a_{j+1}},$ (79)
$=\frac{1}{\pi}\sum_{j}\left(\frac{b_{j}}{a_{j}}+\frac{b_{j+1}}{a_{j+1}}\right),$ (80)
where $b_{j},b_{j+1}\in\mathbb{Z}$ must be chosen to satisfy $b_{j}a_{j+1}+b_{j+1}a_{j}=\tilde{\theta}_{j+\frac{1}{2}}$. Since the $a$ are coprime, we state two consequences of Bezout’s identity [171]. Firstly, $b_{j}$ and $b_{j+1}$ can be chosen for any $\tilde{\theta}_{j+\frac{1}{2}}$. Secondly, $\tilde{\theta}_{j+\frac{1}{2}}$ can be chosen such that $b_{j}$ takes any value mod $a_{j}$, and $b_{j+1}$ takes any value mod $a_{j+1}$ iiiiiiTo illustrate why the latter holds, consider the effect on $b_{j}$ and $b_{j+1}$ when $A_{GS}\rightarrow A_{GS}+a_{j}$..
Consequently, $A_{GS}$ for the whole system can take any value that may be written as
$A_{GS}=w+\sum_{j}\frac{b_{j}}{a_{j}},$ (81)
with $w$ and all $b_{j}$ arbitrarily chosen integers. Thus, there are a total of $\prod_{j}a_{j}$ unique values of $A_{GS}$ between 0 and 1, and thus on any integer-length interval $[x,x+1)$. Note that this product is over all unique values of $a_{j}$ in the system.
Finally, since two configurations are gauge-equivalent only when they differ by a multiple of $u$, the number of degenerate ground states of the composite system is
$N_{GS}=u\prod_{j}a_{j},$ (82)
where the product is again over all unique values of $a_{j}$ present in the system.
@@ IV.1 Ground State Degeneracy
Figure 6: Transport of anyons through a system of fractionalised wires. Each wire may be differently fractionalised with $ua_{j}$, indicated by colour, as well as being unevenly displaced from one another. (a) The $\chi$ operator, Eq. 90, is responsible for transport in the $y$-direction. On a set of three coupled wires, the $\chi_{2}(x)$ operator destroys a composite of $a_{1}$ quasiparticles on the $\tilde{\theta}_{1.5}$ quasiwire, and creates a composite of $a_{3}$ quasiparticles from the $\tilde{\theta}_{2.5}$ quasiwire. (b) The $\rho$ operator, Eq. (95), is responsible for transport in the $x$-direction, where the operator $\rho(x_{1},x_{2})$ destroys a single kink at position $x_{1}$ and creates one at $x_{2}$ on the same wire. (c) On a given $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire, a composite excitation with $a_{j}a_{j+1}$ multiplicity can travel freely to the quasiwire above and below by applying the $\chi_{j}$ operator $a_{j}$ times. Furthermore, the nature of the $\chi$ operator is such that it ensures the new quasiparticle is always created with the correct multiplicity to allow it to travel onto the wire beyond.
IV.2 Quasiparticle Charge
We have seen that the ground state degeneracy of the system of coupled wires has a non-trivial dependence on the fractionalisation of all constituent wires in the system. Since the ground state degeneracy is related to the nature of the quasiparticle excitations, let us now explicitly derive the properties of these excitations.
@@ IV.2 Quasiparticle Charge
Quasiparticles in the bosonised system correspond to a kink in $\tilde{\theta}_{j+\frac{1}{2}}$, where the value jumps by ±$\pm\pi$. As a warm-up, let us start by determining the bare charge carried by quasiparticles in the well-studied uniform case (all $a_{j}=1$), and then move on to the more involved non-uniform case.
@@
The Uniform Case: Following Eq. 25, we may write down the expression for the total bare fermion charge contained on the $j^{\textup{th}}$ wire within a region $x_{I}\rightarrow x_{F}$ as
@@
Using Eqs. 27, 26, 34 and 35 we may now re-express the bare density in terms of our fractionalised density variables $\tilde{\varphi}$ and $\tilde{\theta}$,
@@
$\sum_{j=1}^{N}\frac{\theta_{j}(x)}{\pi}&=\frac{1}{u}\sum_{j=1}^{N-1}\frac{\tilde{\theta}_{j+\frac{1}{2}}}{\pi}+\left(\textup{boundary terms}\right),\end{aligned}$ (86)
@@
The Non-uniform Case: We may once again re-express the bare fermion density in terms of fractionalised fields, although we now must use Eqs. 54 and 55, to get
@@
By considering the denominators in this expression, we see that there are no neat cancellations in $\sum_{j}{\theta_{j}}$, which now retains dependence both on the $\tilde{\varphi}$ and $\tilde{\theta}$ fields in the bulk. This is unsurprising, since we have already established that bare fermionic charge is not conserved in the composite system, see Section III.2.1, unless a substrate is included, see Appendix B. Thus, we should not be surprised to see that overall charge does not have a well-defined expectation value. Rather, let us consider the quantity
@@
IV.3 Composite Statistics
In this section, we calculate anyonic braiding statistics. This calculation takes three parts. In the first we establish the operators necessary to transport anyons throughout the system. Next, we see how the properties of these transport operators lead to the emergence of two types of excitation: a set of composite anyons that are free to travel throughout the system along both axes, and a set of lineons which are confined to move only in the $x$-direction [173, 174]. Finally, we braid both sets of anyons and calculate relative statistics.
IV.3.1 Moving Anyons
@@
[…]
@@
where $\chi$ acts at $x^{\prime}$ and $\partial\tilde{\theta}$ acts at $x$. Thus, we see see that the operator removes $a_{j-1}$ quasiparticles on the $\tilde{\theta}_{j-\frac{1}{2}}$ quasiwire and creates $a_{j+1}$ quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire. A diagram is shown in Fig. 6a. Thus, we see that in the $y$-direction, quasiparticles can only be moved in groups of $a_{j}$.
@@
which is a local operator since $\partial_{x}\tilde{\varphi}$ can always be written in terms of the bare fermion densities, see Eq. 15. This operator allows us to transport a single ±$\pm\pi$ kink along a wire to any arbitrary position.
IV.3.2 Lineons and Composite Anyons
@@
Figure 7: Three types of fractionalised excitations in a system of unevenly spaced wires, where the $j$th wire is in a $1/ua_{j}^{2}$ phase. (a) A single kink between the $j$ and $j+1$ wire corresponds to an excitation carrying ${1}/{ua_{j}^{2}a_{j+1}^{2}}$ statistics, and is constrained to only move in the $x$ direction. (b) A composite of $a_{j}$ kinks between the $j$ and $j+1$ wire is only able to travel up to the next wire and back. (c) Finally, a $C$-anyon, a composite of $a_{k}a_{k+1}$ kinks has complete freedom to travel throughout the system. Note that (a) and (b) are fractons, constrained to move in the $x$-direction, whereas (c) is not constrained.
@@
Following the form of the above quasiparticle transport operators Eqs. 90 and 95, we see that the $x$ and $y$-directions in our model effectively ‘see’ a different type of quasiparticle. In the $x$-direction, each quasiparticle is free to move independently along the wire. However, in the $y$-direction quasiparticles can only be transported together in groups of $a_{j}$, where each interface between wires with a different value of $a_{j}$ thus transfers a different number of quasiparticles, as shown in Fig. 6. Thus, it is natural to ask if there is a minimal combined quasiparticle that is free to travel throughout the system in both directions.
@@
To answer this, we consider a set of $\eta$ quasiparticles at the same position on a quasiwire at $\tilde{\theta}_{j+\frac{1}{2}}$. In order to be able to travel freely in the $y$-direction, we must at least be able to hop this set of quasiparticles both onto the quasiwire above and the one below. Thus, there are two $\chi$ operators to consider, following Eqs. 91 and 92. To move the bundle up to the next wire, we must apply a $\chi_{j+1}$ operator, which destroys $a_{j}$ kinks and creates $a_{j+2}$ kinks on the next quasiwire. Similarly, to move the bundle down, we must apply a $\chi_{j}^{\dagger}$ operator, which destroys $a_{j+1}$ kinks and creates $a_{j-1}$ on the quasiwire below. Thus, for this bundle of quasiparticles to be freely mobile, we must have started with a number $\eta$ of particles in the bundle that is both divisible by $a_{j}$ and $a_{j+1}$. Since all $a_{j}$ are coprime, the minimal acceptable bundle contains $a_{j}\times a_{j+1}$ quasiparticles.
@@
Now, suppose we start with such a bundle of $a_{j}\times a_{j+1}$ quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire and wish to transport it to the quasiwire above. Since $\chi_{j+1}$ only moves $a_{j}$ individual quasiparticles, we will have to apply it $a_{j+1}$ times. Each time we destroy $a_{j}$ particles and create $a_{j+2}$ particles on the next wire. Thus, we finish having removed the full bundle, and created a bundle of $a_{j+1}\times a_{j+2}$ quasiparticles on the quasiwire above, at $\tilde{\theta}_{j+1,j+2}$. Remarkably, the transport operators conspire to ensure that the bundle created on the next wire has the right multiplicity to allow it to hop onto the wire above, and so on. A diagram of this process is shown in Fig. 6c.
@@
Excitations in the system can thus be classified broadly into three types. One has a set of fractionalised quasiparticles corresponding to a single ±$\pm\pi$ kink on a given quasiwire. These are constrained to only move in the $x$-direction – constituting a set of fractonic excitations called lineons [173, 174]. Next, we may combine a group of $a_{j}$ such excitations to form a constrained compound particle, which can only hop over one adjacent wire, but no further. These have also constrained motion as lineons, but are spread among two wires, and hence we refer to them as spread-lineons, or s-lineons for short. Finally, combining $a_{j}\times a_{j+1}$ fractonic quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire produces a free composite quasiparticle that may travel unimpeded throughout the system. We shall label these as $C$-particles. These three cases are shown in Fig. 7.
IV.3.3 Braiding

### v4

@@ IV A General Construction for Non-Uniform Fractionalisation
Figure 6: Transport of anyons through a system of fractionalised wires. Each wire may be differently fractionalised with $ua_{j}$, indicated by colour, as well as being unevenly displaced from one another. (a) The $\chi$ operator, Eq.˜66, is responsible for transport in the $y$-direction. On a set of three coupled wires, the $\chi_{2}(x)$ operator destroys a composite of $a_{1}$ quasiparticles on the $\tilde{\theta}_{1.5}$ quasiwire, and creates a composite of $a_{3}$ quasiparticles from the $\tilde{\theta}_{2.5}$ quasiwire. (b) The $\rho$ operator, Eq. (71), is responsible for transport in the $x$-direction, where the operator $\rho(x_{1},x_{2})$ destroys a single kink at position $x_{1}$ and creates one at $x_{2}$ on the same wire. (c) On a given $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire, a composite excitation with $a_{j}a_{j+1}$ multiplicity can travel freely to the quasiwire above and below by applying the $\chi_{j}$ operator $a_{j}$ times. Furthermore, the nature of the $\chi$ operator is such that it ensures the new quasiparticle is always created with the correct multiplicity to allow it to travel onto the wire beyond.
IV.1 Quasiparticle Charge
@@
Quasiparticles in the bosonised system correspond to a kink in $\tilde{\theta}_{j+\frac{1}{2}}$, where the value jumps by $\pm\pi$. As a warm-up, let us start by determining the bare charge carried by quasiparticles in the well-studied uniform case (all $a_{j}=1$), and then move on to the more involved non-uniform case.
@@
The Uniform Case: Following Eq.˜25, we may write down the expression for the total bare fermion charge contained on the $j^{\textup{th}}$ wire within a region $x_{I}\rightarrow x_{F}$ as
@@
Using Eqs.˜27, 26, 34 and 35 we may now re-express the bare density in terms of our fractionalised density variables $\tilde{\varphi}$ and $\tilde{\theta}$,
@@
$\begin{aligned} \sum_{j=1}^{N}\frac{\theta_{j}(x)}{\pi}&=\frac{1}{u}\sum_{j=1}^{N-1}\frac{\tilde{\theta}_{j+\frac{1}{2}}}{\pi}+\left(\textup{boundary terms}\right),\end{aligned}$ (62)
@@
The Non-uniform Case: We may once again re-express the bare fermion density in terms of fractionalised fields, although we now must use Eqs.˜54 and 55, to get
@@
By considering the denominators in this expression, we see that there are no neat cancellations in $\sum_{j}{\theta_{j}}$, which now retains dependence both on the $\tilde{\varphi}$ and $\tilde{\theta}$ fields in the bulk. This is unsurprising, since we have already established that bare fermionic charge is not conserved in the composite system, see Section˜III.2.1, unless a substrate is included, see Appendix˜B. Thus, we should not be surprised to see that overall charge does not have a well-defined expectation value. Rather, let us consider the quantity
@@
IV.2 Composite Statistics
In this section, we calculate anyonic braiding statistics. This calculation takes three parts. In the first we establish the operators necessary to transport anyons throughout the system. Next, we see how the properties of these transport operators lead to the emergence of two types of excitation: a set of composite anyons that are free to travel throughout the system along both axes, and a set of lineons which are confined to move only in the $x$-direction [169, 170, 176]. Finally, we braid both sets of anyons and calculate relative statistics.
IV.2.1 Moving Anyons
@@
$\begin{aligned} \chi_{j}(x)=&{\psi^{L}_{j}}^{\dagger}(x){\psi^{R}_{j}}^{\phantom{}}(x)\\ &\rightarrow e^{-i[\phi^{R}_{j}(x)-\phi^{L}_{j}(x)]}.\end{aligned}$ (65)
@@
where $\chi$ acts at $x^{\prime}$ and $\partial\tilde{\theta}$ acts at $x$. Thus, we see see that the operator removes $a_{j-1}$ quasiparticles on the $\tilde{\theta}_{j-\frac{1}{2}}$ quasiwire and creates $a_{j+1}$ quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire. A diagram is shown in Fig.˜6a. Thus, we see that in the $y$-direction, quasiparticles can only be moved in groups of $a_{j}$.
@@
which is a local operator since $\partial_{x}\tilde{\varphi}$ can always be written in terms of the bare fermion densities, see Eq.˜15. This operator allows us to transport a single $\pm\pi$ kink along a wire to any arbitrary position.
IV.2.2 Lineons and Composite Anyons
@@
Figure 7: Three types of fractionalised excitations in a system of unevenly spaced wires, where the $j$th wire is in a $1/ua_{j}^{2}$ phase. (a) A single kink between the $j$ and $j+1$ wire corresponds to a quasiparticle excitation that is constrained to only move in the $x$ direction. (b) A composite of $a_{j}$ kinks between the $j$ and $j+1$ wire is only able to travel up to the next wire and back. (c) Finally, a $C$-anyon, a composite of $a_{k}a_{k+1}$ kinks has complete freedom to travel throughout the system. Note that (a) and (b) are fractons, constrained to move in the $x$-direction, whereas (c) is not constrained.
@@
Following the form of the above quasiparticle transport operators Eqs.˜66 and 71, we see that the $x$ and $y$-directions in our model effectively ‘see’ a different type of quasiparticle. In the $x$-direction, each quasiparticle is free to move independently along the wire. However, in the $y$-direction quasiparticles can only be transported together in groups of $a_{j}$, where each interface between wires with a different value of $a_{j}$ thus transfers a different number of quasiparticles, as shown in Fig.˜6. Thus, it is natural to ask if there is a minimal combined quasiparticle that is free to travel throughout the system in both directions.
@@
To answer this, we consider a set of $\eta$ quasiparticles at the same position on a quasiwire at $\tilde{\theta}_{j+\frac{1}{2}}$. In order to be able to travel freely in the $y$-direction, we must at least be able to hop this set of quasiparticles both onto the quasiwire above and the one below. Thus, there are two $\chi$ operators to consider, following Eqs.˜67 and 68. To move the bundle up to the next wire, we must apply a $\chi_{j+1}$ operator, which destroys $a_{j}$ kinks and creates $a_{j+2}$ kinks on the next quasiwire. Similarly, to move the bundle down, we must apply a $\chi_{j}^{\dagger}$ operator, which destroys $a_{j+1}$ kinks and creates $a_{j-1}$ on the quasiwire below. Thus, for this bundle of quasiparticles to be freely mobile, we must have started with a number $\eta$ of particles in the bundle that is both divisible by $a_{j}$ and $a_{j+1}$. Since all $a_{j}$ are coprime, the minimal acceptable bundle contains $a_{j}\times a_{j+1}$ quasiparticles.
@@
Now, suppose we start with such a bundle of $a_{j}\times a_{j+1}$ quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire and wish to transport it to the quasiwire above. Since $\chi_{j+1}$ only moves $a_{j}$ individual quasiparticles, we will have to apply it $a_{j+1}$ times. Each time we destroy $a_{j}$ particles and create $a_{j+2}$ particles on the next wire. Thus, we finish having removed the full bundle, and created a bundle of $a_{j+1}\times a_{j+2}$ quasiparticles on the quasiwire above, at $\tilde{\theta}_{j+1,j+2}$. Remarkably, the transport operators conspire to ensure that the bundle created on the next wire has the right multiplicity to allow it to hop onto the wire above, and so on. A diagram of this process is shown in Fig.˜6c.
@@
Excitations in the system can thus be classified broadly into three types. One has a set of fractionalised quasiparticles corresponding to a single $\pm\pi$ kink on a given quasiwire. These are constrained to only move in the $x$-direction – constituting a set of fractonic excitations called lineons [169, 170]. Next, we may combine a group of $a_{j}$ such excitations to form a constrained compound particle, which can only hop over one adjacent wire, but no further. These have also constrained motion as lineons, but are spread among two wires, and hence we refer to them as spread-lineons, or s-lineons for short. Finally, combining $a_{j}\times a_{j+1}$ fractonic quasiparticles on the $\tilde{\theta}_{j+\frac{1}{2}}$ quasiwire produces a free composite quasiparticle that may travel unimpeded throughout the system. We shall label these as $C$-particles. These three cases are shown in Fig.˜7.
IV.2.3 Braiding